OpenStudy (anonymous):

A particular model of walkie talkie can broadcast in a circular area. The radius of the broadcast area is 10,000 feet. Find the area of this circle to the hearest square foot.

OpenStudy (anonymous):

\[ A = \pi r^2 = 3.14 159 265 \times 10000^2 = 314 159 265 \]